OpenStudy (anonymous):

fred begins walking toward johns house at 3mi/h. john leaves his house at the same time and walks towards fred house on the same path at a rate of 2mi/h. How long will it be before they meet if the distance between their houses is 4mi?

OpenStudy (anonymous):

ok got it

OpenStudy (anonymous):

so this is what it looks like right

OpenStudy (anonymous):

k

OpenStudy (anonymous):

wait like thats how u set it up or like how u write it

OpenStudy (anonymous):

ok lets start over fred=a john=b

OpenStudy (anonymous):

and forget about the drawing

OpenStudy (anonymous):

alright

OpenStudy (anonymous):

if fred walks 3mi/hr and john walks 2mi/hr then they walk a total of 5mi/hr correct

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

so we can use the variable t for time so 3t+2t=4 correct

OpenStudy (anonymous):

4t

OpenStudy (anonymous):

wait no it equals 5t

OpenStudy (anonymous):

yes so solve for t and what do you get?

OpenStudy (anonymous):

solve 3t+2t=5t

OpenStudy (anonymous):

so 5t=4 solve

OpenStudy (anonymous):

t=1.25

OpenStudy (anonymous):

if you divide both sides by 5 you get t=4/5

OpenStudy (anonymous):

dont u divide 4 by 5. like 5 divided by 4

OpenStudy (anonymous):

nope 4/5

OpenStudy (anonymous):

how

OpenStudy (anonymous):

lol idk how its just the way math works

OpenStudy (anonymous):

its wierd though cause 5 has the variable. wouldnt u divide both sides by 5

OpenStudy (anonymous):

yes you do divide both sides by 5 so you get t=4/5

OpenStudy (anonymous):

oh i get it

OpenStudy (anonymous):

lol

OpenStudy (anonymous):

ya so now that you have 4/5hr. you need to convert that into min. do you know how to do that?

OpenStudy (anonymous):

omg this is so much work. umm..sec into min, min into hours

OpenStudy (anonymous):

really easy actually just multiple 60*(4/5)

OpenStudy (anonymous):

i put 60 over 1?

OpenStudy (anonymous):

yes and multiply across

OpenStudy (anonymous):

k. so it will be 1200

OpenStudy (anonymous):

no try this multiply 60 by 4 60*4

OpenStudy (anonymous):

240. god im really kinda confused right now

OpenStudy (anonymous):

ok now divide 240 by 5 240/5

OpenStudy (anonymous):

u dont cross multiply because like the one ontop are the same thing so u have to multiply those together?? if it makes sense

OpenStudy (anonymous):

48

OpenStudy (anonymous):

yes thats your answer

OpenStudy (anonymous):

dont u multiply240 by 5

OpenStudy (anonymous):

oh i get it

OpenStudy (anonymous):

no think of 4/5 as .8

OpenStudy (anonymous):

uuhh now im really confused. i think imma ask my teacher tomorrow. I think dis is a problem that i need to see to understand. Thank u soo much though :D

OpenStudy (anonymous):

im sorry dont know how to explain it any differently

OpenStudy (anonymous):

this is what we did though Let t=time spent walking until they meet distance(d)=rate(r) times time(t) or d=rt;r=d/t and t=d/r distance John walked=rt=3*t distance Fred walked =rt=2*t Now we know that when they meet, the distance John had walked plus the distance Fred had walked must equal 4 mi. So our equation to solve is: 3t+2t=4 5t=4 divide both sides by 5 t=4/5 hr or 48 min-------------time elapsed until they meet CK 3*(4/5)+2*(4/5)=4 mi 12/5+8/5=4 20/5=4 4=4